ACCIDENTAL CONICAL INTERSECTIONS IN MIXED TRIMERS OF POTASSIUM AND RUBIDIUM: A VIBRONIC ANALYSIS OF THE 4^{4}B_{2} AND 3^{4}A_{1} STATES

\underline{A. W. HAUSER}, G. AUBÖCK, C. CALLEGARI^{a} and W. E. ERNST, Institute of Experimental Physics, Graz University of Technology, Petersgasse 16, A-8010 Graz, Austria.

We compare the 3^{4}A_{1} and 4^{4}B_{2} states of homonuclear and heteronuclear alkali trimers formed of potassium and rubidium. The Multireference Rayleigh Schrödinger Perturbation Theory of second order is applied to obtain the corresponding adiabatic potential energy surfaces. In the case of homonuclear trimers these pairs of states correspond to the two branches of the E×e Jahn-Teller distorted 2^{4}E′ state. For heteronuclear trimers, the vibrational modes Q_{x} and Q_{y} are no longer degenerate, but the two electronic states still show a conical intersection at obtuse (KRb_{2}) or acute (K_{2}Rb) isosceles geometries. Spectroscopic consequences of this situation are discussed, vibronic spectra are predicted and compared to LIF spectra obtained in helium droplet isolation spectroscopy experiments of our group^{b}^{,}^{c}.

apresent adress: Sincrotrone Trieste, Strada Statale 14 - km 163.5, 34149 Basovizza, Trieste, Italy

bJ. Nagl, G. Auböck, A.W. Hauser, O. Allard, C. Callegari and W.E. Ernst, Phys. Rev. Lett. 100, 063001 (2008)

cJ. Nagl, G. Auböck, A.W. Hauser, O. Allard, C. Callegari and W.E. Ernst, J. Chem. Phys. 128, 154320 (2008)